# Implementing Evidence Acquisition:

### v2

@@ 3.2 A Set of Static Problems from the Dynamic Problem
We now introduce key tools that will be useful in proving our results. Specifically, we decompose the dynamic problem into a sequence of static problems. Each static game in our decomposition is indexed by a pair of times, $t,t^{\prime}\in[0,T]$, which we refer to as the continuation game between $t$ and $t^{\prime}$, denoted $\mathcal{G}_{t,t^{\prime}}$. Each continuation game considers the agent’s problem at time $t$, but drops all incentive constraints except for (a) the one at time $t$, which says the agent is willing to start working, and (b) the one at $t^{\prime}$, which says they are willing to report the truth after having followed the stopping strategy which stops at $t^{\prime}$.
@@
In the original dynamic environment, the agent has incentives to exert effort at time $t$ if there exists $t^{\prime}\geq t$ such that the agent has incentives to exert effort in the continuation game $\mathcal{G}_{t,t^{\prime}}$. If the agent has incentives to work in every continuation game $\mathcal{G}_{t}$, then by backward induction the contract can implement $\tau_{R}$. The natural question is whether effort is implementable in a given continuation game, a question whose answer follows immediately from existing results for the static case. To describe this we use the following concept:
@@
A scoring rule $P$ is V-shaped with parameters $z_{0},z_{1}$ if
$P(\mu,\theta)=\begin{cases}z_{1}&\mu\geq\frac{z_{0}}{z_{1}+z_{0}}\text{ and }\theta=1\\ z_{0}&\mu<\frac{z_{0}}{z_{1}+z_{0}}\text{ and }\theta=0\\ 0&\text{otherwise}.\end{cases}$
We say $P$ is a V-shaped scoring rule with kink at $D$ if the parameters $z_{0},z_{1}$ satisfies $z_{0}=1,z_{1}=\frac{1-D}{D}$ if $D\geq\frac{1}{2}$ and $z_{0}=\frac{D}{1-D},z_{1}=1$ if $D<\frac{1}{2}$.
@@
Proposition 1 illustrates the tensions involved in designing dynamic contracts. As the agent’s posterior evolves over time, the priors for the continuation games at different time periods vary, leading to inconsistencies in the scoring rules that maximize incentives for effort across these periods. In particular, as illustrated in Section 3.1, to implement maximum effort in dynamic environments, the moral hazard constraints bind at both time 00 and the stopping time $\tau$ when the signals are perfectly revealing. A V-shaped scoring rule with a kink at $\mu^{N}_{\tau}$ leads to insufficient incentives for the agent to exert effort at time 00, resulting in the agent not starting work at all. Conversely, a V-shaped scoring rule with a kink at $\mu^{N}_{0}$ results in insufficient incentives for the agent to exert effort at time $\tau$, causing the agent to stop prematurely. Our illustration demonstrates that to balance the incentives for exerting effort across different time periods, the V-shaped scoring rule may need a kink located at some interior belief. Moreover, as we will show later in Section 4.3, the optimal scoring rule may not take a V-shaped form to provide balanced incentives in dynamic environments.
@@ 4.1 Stationary Environment
The simplest case of interest is the stationary environment, where the agent’s no information belief is $\mu^{N}_{t}=D$ for all $t\leq T$. In this simple case, all continuation games at any time $t$ share the same prior belief. But Proposition 1 showed that the prior determines the effort-maximizing scoring rule for all continuation games. Thus, it is immediate that the V-shaped scoring rule with kink at the prior $D$ can implement any implementable effort level.
@@ 4.2 Perfect-learning Environment
Outside of the stationary environment, the no-information belief $\mu^{N}_{t}$ of the agent drifts over time, so the aforementioned tension in designing effort-maximizing scoring rules does arise: In particular, it could be that there exist scoring rules which can implement effort from 0 to $\tilde{t}$ and from $\tilde{t}$ to $\tau$, but no dynamic contract that can implement effort from 0 to $\tau$. This idea was mentioned in Section 3.1, where we argued that (a) a scoring rule with $r_{1}<1$ might be necessary to induce the agent to exert effort from some $\mu_{0}^{N}>\mu^{*}$ to $\mu_{t}^{N}=1/2$ (for instance), while (b) a scoring rule with $r_{1}=1$ would be necessary to induce the agent to exert effort from $\mu_{t}^{N}=1/2$ to $\mu_{\tau}^{N}=\frac{c}{\lambda_{1}^{G}}$, but (c) no contract (static or dynamic) could induce the agent to exert effort between $\mu_{0}^{N}$ and $\frac{c}{\lambda_{1}^{G}}$. Thus, adjusting the solutions to the relaxed problems is necessary to find a scoring rule that is also an effort-maximizing contract.
We now show that any implementable effort level can be implemented via a scoring rule under perfect-learning. Unlike for the stationary environment, while a V-shaped scoring rule implements maximum effort, its kink need not be at the prior $D$:
@@
In the perfect-learning environment, a V-shaped scoring rule with kink at $\frac{1}{1+r_{1}}\in(1/2,1)$ implements maximum effort.
@@
The proof of Theorem 2 does not identify an effort-maximizing choice of $r_{1}$, but rather argues that any dynamic contract (within the class identified in Lemma 3) can be replaced by a scoring rule while increasing the incentives to exert effort. We provide some additional results to show how to identify the effort-maximizing choice of $r_{1}$ and, in particular, show why this kink need not be at the prior (in contrast the static case). This step also provides formal details behind an interesting dynamic effect in our model alluded to in Section 3.1: for long time horizons, there exist $\mu_{1}>\mu_{2}>\mu_{3}$ such that the agent can be incentivized to work when the no information belief would drift from (a) $\mu_{1}$ to $\mu_{2}$ or (b) $\mu_{2}$ to $\mu_{3}$, but not when this belief would drift from $\mu_{1}$ to $\mu_{3}$.
@@ 4.3 Single-signal Environment
The correct intuition is as follows: while providing a strictly positive base reward at time $t$ decreases the agent’s incentive to exert effort at time $t$, it increases the agent’s incentive to exert effort at earlier times $t^{\prime}<t$; indeed, the agent may expect high rewards for exerting effort, even if mistakes are made in guessing a state after receiving imperfect signals. This modification induces more effort if the agent’s incentive constraint for exerting effort initially binds at time 0 but becomes slack at intermediate time $t\in(0,\tau_{R})$. As illustrated in Figure 4 and in Section 4.2, by implementing the effort-maximizing V-shaped scoring rule, the agent’s incentive for exerting effort is binding only at the extreme time 00 with belief $D=\bar{\mu}(r_{1})$ and time $\tau_{R}$ with belief $\mu^{N}_{\tau_{R}}=\underline{\mu}(r_{1})$. In this case, since the signals are not perfectly revealing, there may exist a time $t$ such that $\mu^{G}_{t}<D$. By providing an additional menu option with a strictly positive base reward in the scoring rule to increase the agent’s utility at beliefs $\mu^{G}_{t}$ (e.g., the additional menu option $(z_{0},z_{1})$ illustrated in Figure 4), the agent’s incentive constraint for exerting effort at time 00 is relaxed and the contract thus provides the agent incentives to exert effort following more extreme prior beliefs without influencing the stopping belief $\underline{\mu}(r_{1})$.
@@
While convexity of the no-information utility function does imply that the effort-maximizing contract has an implementation as a scoring rule—since convex functions are equal to the upper envelope of the linear functions below them—the resulting scoring rule need not satisfy the reward constraint, i.e., that $r_{\theta}\in[0,1]$. For instance, consider a convex no-information utility function $u^{N}_{t}=(\mu^{N}_{t})^{2}$. A simple dynamic contract that implements this no-information utility function is to offer a constant reward $(\mu^{N}_{t})^{2}$ at time $t$ regardless of the realization of the state. However, to implement this utility function using a scoring rule, by Lemma 2, the menu option for belief $\mu\in[0,1]$ must be $(-\mu^{2},2\mu-\mu^{2})$, which violates the ex post individual rationality constraint.
@@ 5.1 Conditions for the Insufficiency of Scoring Rule
This section shows that the effort-maximizing scoring rule must be dynamic when (i) signals are noisy and (ii) the drift is sufficiently slow. Specifically, we consider cases where both good and bad news signals can arrive with strictly positive probability if the agent exerts effort, so that $\lambda_{1}^{G}>\lambda_{0}^{G}>0$ and $0<\lambda_{1}^{B}<\lambda_{0}^{B}$. Recall that we assumed without loss that $\lambda_{1}^{G}+\lambda_{1}^{B}>\lambda_{0}^{G}+\lambda_{0}^{B}$, so beliefs drift toward state 0 absent news. Taking the drift to be slow means that $\lambda_{1}^{G}+\lambda_{1}^{B}$ is only slightly larger than $\lambda_{0}^{G}+\lambda_{0}^{B}$. Signals being noisy means that they do not reveal the state.
@@
and $T\geq T_{\lambda,D,c}$, any static scoring rule implements effort strictly less than the maximum.
We discuss the role of each of the four conditions on $\lambda$. Two essentially avoid degenerate cases. Specifically, the condition $T\geq T_{\lambda,D,c}$ implies the time horizon $T$ will not be a binding constraint for the agent to exert effort. The sufficient incentive condition avoids trivial solutions by ensuring that the agent has incentives to exert effort for a strictly positive length of time in the effort-maximizing contract. More substantive in light of our previous results, however, are the other two conditions: The noisy-signal condition rules out the perfect-learning environment and the single-signal environment, and the slow-drift condition rules out the stationary environment. Moreover, the single-signal environment can also be viewed as the extreme opposite of the slow-drift condition, since the difference in arrival rates of the signals is maximal and the belief drifts to 0 quickly in the absence of a Poisson arrival.
A substantive assumption, but one that appears relatively harmless, is that $D\in(0,\frac{1}{2})$. Note that combining this with our restriction to environments where the belief absent a Poisson signal drifts towards 0, our theorem only applies to environments where the agent’s posterior belief absent a Poisson signal becomes more polarized towards the more likely states given the prior belief. This restriction ensures that the following myopic-incentive contract we define is incentive compatible for the agent with rewards in $[0,1]$.
@@
By contrast, if the agent instead faced a contract that can be implemented as a scoring rule, given stopping time $\tau_{R^{*}}$ in the effort-maximizing contract $R^{*}$, this scoring rule must be close to the effort-maximizing scoring rule in the continuation game $\mathcal{G}_{t}$ for any $t$ that is sufficiently close to $\tau_{R^{*}}$ in order to incentivize the agent to exert effort in continuation game $\mathcal{G}_{t}$. However, when the signals are noisy, as illustrated in Appendix OA 4, such scoring rules fail to provide sufficient incentives for the agent to exert effort at time 00, leading to a contradiction. Intuitively, the myopic-incentive contract avoids this conflict in incentives by providing higher rewards to the agent upon receiving a bad new signal $B$ without affecting the agent’s incentive to report the acquired information truthfully. In particular, higher rewards upon Poisson signal arrivals strengthen the agent’s incentives to exert effort at any time $t$. Interestingly, when $\varepsilon=0$, the environment is stationary, in which case a scoring rule is again optimal; however, as long as drift is non-zero, scoring rules will fail to approximate the optimum in contrast to the myopic-incentive contracts.
5.2 Effort-Maximizing Dynamic Contracts
Lemma 7 shows that the myopic-incentive contract approximately implements maximum effort when the drift is slow and the belief absent a Poisson signal is polarizing. Although the stated contract is generally not fully effort-maximizing, a similar decreasing reward structure does, in fact, characterize effort-maximizing mechanisms.
@@ 5.2 Effort-Maximizing Dynamic Contracts
The only feature of myopic-incentive contracts that does not extend to effort-maximizing contracts is the rate of decrease for rewards following “bad news” signals. In myopic-incentive contracts, rewards for “bad news” signals are chosen such that the menu options $(r^{G}_{t},r^{B}_{t})$ offered at time $t$ consist of a V-shaped scoring rule with a kink at belief $\mu^{N}_{t}$. Such a rate of decrease is shown to approximately implement maximum effort (Lemma 7), but it implements strictly less than the maximum in general. In general environments, the effort-maximizing rate of decreasing rewards for bad news signals depends on primitives and may not admit closed-form characterizations. However, such rewards can be computed efficiently by solving a family of linear programs (Appendix OA 2).
@@
6 Extensions and Discussions
We now turn to some extensions of the main model, particularly those related to randomized contracts, non-stationary information acquisition, and ex-ante reward constraints. We show how to generalize our results and techniques to such environments. Finally, we conclude our paper and propose several open questions in Section 7.

### v3

@@ 3.2 A Set of Static Problems from the Dynamic Problem
We now introduce key tools that will be useful in proving our results. Specifically, we decompose the dynamic problem into a sequence of static problems. Each static game in our decomposition is indexed by a pair of times, $t,t^{\prime}\in[\Delta,T]$, which we refer to as the continuation game between $t$ and $t^{\prime}$, denoted $\mathcal{G}_{t,t^{\prime}}$. Each continuation game considers the agent’s problem at time $t$, but drops all incentive constraints except for (a) the one at time $t$, which says the agent is willing to start working, and (b) the one at $t^{\prime}$, which says they are willing to report the truth after having followed the stopping strategy which stops at $t^{\prime}$.
@@
In the original dynamic environment, at any time $t$, the agent has incentives to exert effort at time $t\leq\tau$ if and only his continuation payoff $V_{t}(\mu^{N}_{t})$ is at least his payoff from not exerting effort. This is equivalent to the condition that the agent has incentives to exert effort in the continuation game $\mathcal{G}_{t}$. Therefore, before investigating the optimal dynamic contract, a natural question is whether effort is implementable in a given continuation game. The answer to this question follows directly from existing results for the static case. To formalize this, we introduce the following concept:
@@
A scoring rule $P$ is V-shaped with parameters $r_{0},r_{1}$ if
$P(\mu,\theta)=\begin{cases}r_{1}&\mu\geq\frac{r_{0}}{r_{1}+r_{0}}\text{ and }\theta=1\\ r_{0}&\mu<\frac{r_{0}}{r_{1}+r_{0}}\text{ and }\theta=0\\ 0&\text{otherwise}.\end{cases}$
We say $P$ is a V-shaped scoring rule with kink at $\hat{\mu}\in[0,1]$ if the parameters $r_{0},r_{1}$ satisfies $r_{0}=1,r_{1}=\frac{1-\hat{\mu}}{\hat{\mu}}$ if $\hat{\mu}\geq\frac{1}{2}$ and $r_{0}=\frac{\hat{\mu}}{1-\hat{\mu}},r_{1}=1$ if $\hat{\mu}<\frac{1}{2}$.
@@
Proposition 1 illustrates the tensions involved in designing dynamic contracts. As the agent’s posterior evolves over time, the priors for the continuation games at different time periods vary, leading to inconsistencies in the scoring rules that maximize incentives for effort across these periods. In particular, to implement maximum effort in dynamic environments, the moral hazard constraints bind at both time 00 and the stopping time $\tau$ when the signals are perfectly revealing. A V-shaped scoring rule with a kink at $\mu^{N}_{\tau}$ leads to insufficient incentives for the agent to exert effort at time 00, resulting in the agent not starting work at all. Conversely, a V-shaped scoring rule with a kink at $\mu^{N}_{0}$ results in insufficient incentives for the agent to exert effort at time $\tau$, causing the agent to stop prematurely. Our illustration demonstrates that to balance the incentives for exerting effort across different time periods, the V-shaped scoring rule may need a kink located at some interior belief. Moreover, as we will show later in Section 4.3, the optimal scoring rule may not take a V-shaped form to provide balanced incentives in dynamic environments.
@@ 4.1 Stationary Environment
The simplest case of interest is the stationary environment, where the agent’s no information belief is $\mu^{N}_{t}=D$ for all $t\leq T$. In this simple case, all continuation games at any time $t$ share the same prior belief. But Proposition 1 showed that the prior determines the effort-maximizing scoring rule for all continuation games. Thus, it is immediate that the V-shaped scoring rule with kink at the prior $D$ can implement any implementable effort level. [^12] As mentioned in Section 2.1, since Proposition 1 holds for general information acquisition technologies, this result also holds even if $\left\lvert{S}\right\rvert>2$, with an identical proof.
@@ 4.2 Perfect-learning Environment
Outside of the stationary environment, the no-information belief $\mu^{N}_{t}$ of the agent drifts over time, so the aforementioned tension in designing effort-maximizing scoring rules does arise: In particular, it could be that there exist scoring rules which can implement effort from 0 to $\tilde{t}$ and from $\tilde{t}$ to $\tau$, but no dynamic contract that can implement effort from 0 to $\tau$. Adjusting the solutions to the relaxed problems is necessary to find a scoring rule that is also an effort-maximizing contract.
In contrast to the stationary environment, while a V-shaped scoring rule implements maximum effort, its kink need not be at the prior $D$:
@@
In the perfect-learning environment, a V-shaped scoring rule with parameters $r_{0}=1,r_{1}\in(0,1)$ (and hence, kink at $\frac{1}{1+r_{1}}\in(1/2,1)$) implements maximum effort.
@@
4.2.1 Maximizing Effort under Perfect Learning
The proof of Theorem 2 does not identify an effort-maximizing choice of $r_{1}$, but rather argues that any dynamic contract (within the class identified in Lemma 3) can be replaced by a scoring rule while increasing the incentives to exert effort. We provide some additional results to show how to identify the effort-maximizing choice of $r_{1}$ and, in particular, show why this kink need not be at the prior (in contrast the static case). This step also provides formal details behind an interesting dynamic effect in our model: for long time horizons, there exist $\mu_{1}>\mu_{2}>\mu_{3}$ such that the agent can be incentivized to work when the no information belief would drift from (a) $\mu_{1}$ to $\mu_{2}$ or (b) $\mu_{2}$ to $\mu_{3}$, but not when this belief would drift from $\mu_{1}$ to $\mu_{3}$.
@@ 4.3 Single-signal Environment
The correct intuition is as follows: while providing a strictly positive base reward at time $t$ decreases the agent’s incentive to exert effort at time $t$, it increases the agent’s incentive to exert effort at earlier times $t^{\prime}<t$; indeed, the addition of the positive base reward leads the agent to anticipate higher rewards from exerting effort in cases where the terminal belief is in an intermediate range. This modification induces more effort if the agent’s incentive constraint for exerting effort initially binds at time 0 but becomes slack at intermediate time $t\in(0,\tau_{R})$. As illustrated in Figure 5 and in Section 4.2, by implementing the effort-maximizing V-shaped scoring rule, the agent’s incentive for exerting effort is binding only at the extreme time 00 with belief $D=\bar{\mu}(r_{1})$ and time $\tau_{R}$ with belief $\mu^{N}_{\tau_{R}}=\underline{\mu}(r_{1})$. In this case, since the signals are not perfectly revealing, there may exist a time $t$ such that $\mu^{G}_{t}<D$. By providing an additional menu option with a strictly positive base reward in the scoring rule to increase the agent’s utility at beliefs $\mu^{G}_{t}$ (e.g., the additional menu option $(z_{0},z_{1})$ illustrated in Figure 5), the agent’s incentive constraint for exerting effort at time 00 is relaxed and the contract thus provides the agent incentives to exert effort following more extreme prior beliefs without influencing the stopping belief $\underline{\mu}(r_{1})$.
@@
While convexity of the no-information utility function does imply the effort-maximizing contract has an implementation as a scoring rule—since convex functions are equal to the upper envelope of the linear functions below them—the resulting scoring rule need not satisfy the reward constraint, i.e., that $r_{\theta}\in[0,1]$. For instance, consider a convex no-information utility function $u^{N}_{t}=(\mu^{N}_{t})^{2}$. A simple dynamic contract that implements this no-information utility function is to offer a constant reward $(\mu^{N}_{t})^{2}$ at time $t$ regardless of the realization of the state. However, to implement this utility function using a scoring rule, by Lemma 2, the menu option for belief $\mu\in[0,1]$ must be $(-\mu^{2},2\mu-\mu^{2})$, which violates the ex post individual rationality constraint.
@@ 5.1 Conditions for the Insufficiency of Scoring Rule
As alluded to in Section 3.1, contracts eliciting only summarized advice at a fixed date may not implement maximum effort outside of the perfect learning and single-signal environments. Here we show that the effort-maximizing contract must be dynamic when (i) signals are noisy and (ii) the drift is sufficiently slow. Specifically, we consider cases where both good and bad news signals can arrive with strictly positive probability if the agent exerts effort, so that $\lambda_{1}^{G}>\lambda_{0}^{G}>0$ and $0<\lambda_{1}^{B}<\lambda_{0}^{B}$. Recall that we assumed without loss that $\lambda_{1}^{G}+\lambda_{1}^{B}>\lambda_{0}^{G}+\lambda_{0}^{B}$, so beliefs drift toward state 0 absent news. Taking the drift to be slow means that $\lambda_{1}^{G}+\lambda_{1}^{B}$ is only slightly larger than $\lambda_{0}^{G}+\lambda_{0}^{B}$. Signals being noisy means that they do not reveal the state.
@@
  • and $T\geq T_{\lambda,D,c}$, (sufficient-horizon)
, any static scoring rule implements effort strictly less than the maximum.
We discuss the role of each condition. Two essentially avoid degenerate cases. Specifically, the condition $T\geq T_{\lambda,D,c}$ implies the time horizon $T$ will not be a binding constraint for the agent to exert effort. The sufficient incentive condition avoids trivial solutions by ensuring that the agent has incentives to exert effort for a strictly positive length of time in the effort-maximizing contract. More substantive in light of our previous results, however, are the other two conditions: The noisy-signal condition rules out the perfect-learning environment and the single-signal environment, and the slow-drift condition rules out the stationary environment. Moreover, the single-signal environment can also be viewed as the extreme opposite of the slow-drift condition, since the difference in arrival rates of the signals is maximal and the belief drifts to 0 quickly in the absence of a Poisson arrival.
@@
By contrast, if the agent instead faced a contract that can be implemented as a scoring rule, given stopping time $\tau_{R^{*}}$ in the effort-maximizing contract $R^{*}$, this scoring rule must be close to the effort-maximizing scoring rule in the continuation game $\mathcal{G}_{t}$ for any $t$ that is sufficiently close to $\tau_{R^{*}}$ in order to incentivize the agent to exert effort in continuation game $\mathcal{G}_{t}$. However, when the signals are noisy, as illustrated in Appendix OA 3, such scoring rules fail to provide sufficient incentives for the agent to exert effort at time 00, leading to a contradiction. Intuitively, the myopic-incentive contract avoids this conflict in incentives by providing higher rewards to the agent upon receiving a bad new signal $B$ without affecting the agent’s incentive to report the acquired information truthfully. In particular, higher rewards upon Poisson signal arrivals strengthen the agent’s incentives to exert effort at any time $t$. Interestingly, when $\varepsilon=0$, the environment is stationary, in which case a scoring rule is again optimal; however, as long as drift is non-zero, scoring rules will fail to approximate the optimum in contrast to the myopic-incentive contracts.
We mention that substantive assumption in Theorem 4, but one that appears relatively harmless, is that $D\in(0,\frac{1}{2})$. Note that combining this with our restriction to environments where the belief absent a Poisson signal drifts towards 0, our theorem only applies to environments where the agent’s posterior belief absent a Poisson signal becomes more polarized towards the more likely states given the prior belief. This restriction ensures that the following myopic-incentive contract we define is incentive compatible for the agent with rewards in $[0,1]$.
5.2 General Properties of Effort-Maximizing Contracts
Lemma 7 shows that the myopic-incentive contract approximately implements maximum effort when the drift is slow and the belief absent a Poisson signal is polarizing. Although the stated contract is generally not fully effort-maximizing, a similar decreasing reward structure does, in fact, characterize effort-maximizing mechanisms, as illustrated in our numerical exercise in Section 3.1.
@@ 5.2 General Properties of Effort-Maximizing Contracts
The only feature of myopic-incentive contracts that does not extend to effort-maximizing contracts is the rate of decrease for rewards following “bad news” signals. In myopic-incentive contracts, rewards for “bad news” signals are chosen such that the menu options $(r^{G}_{t},r^{B}_{t})$ offered at time $t$ consist of a V-shaped scoring rule with a kink at belief $\mu^{N}_{t}$. Such a rate of decrease is shown to approximately implement maximum effort (Lemma 7), but it implements strictly less than the maximum in general. In general environments, the effort-maximizing rate of decreasing rewards for bad news signals depends on primitives and may not admit closed-form characterizations. But as shown in Section 3.1, such rewards can be computed efficiently by solving a family of linear programs.
@@
6 Additional Observations and Final Thoughts
We end with some additional discussion of our modelling assumptions, illustrating how our results would change under alternative assumptions. We present our concluding thoughts and propose several open questions in Section 6.3.